OpenStudy (anonymous):

What is the mean of the set of numbers? {–9, –8, 4, –1} A. –5.5 B. –3.5 C. –2.8 D. 14 Question Resources

OpenStudy (solomonzelman):

The mean (or the average), is found by adding all the numbers in the set, and dividing by the number of these numbers.

OpenStudy (solomonzelman):

In your case: \(\mu =\dfrac{(-9)+(-8)+(4)+(-1)}{4}\) where \(\mu\) denotes the mean.

OpenStudy (anonymous):

the answer is 4

OpenStudy (solomonzelman):

What is the sum of all elements in your set, can you find that for me?

OpenStudy (solomonzelman):

(-9) + (-8) + (4) + (-1) = ?

OpenStudy (anonymous):

-14

OpenStudy (solomonzelman):

yes, that is right.

OpenStudy (anonymous):

so thats the answer

OpenStudy (solomonzelman):

Now, lets get back to finding the \(\mu\) (the mean). \(\mu =\dfrac{(-9)+(-8)+(4)+(-1)}{4}=\dfrac{-14}{4}=?\)

OpenStudy (anonymous):

-3.5

OpenStudy (solomonzelman):

Yes, and this is exactly right!

OpenStudy (anonymous):

so that is the final answer

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

thank you so much

OpenStudy (solomonzelman):

You are welcome!